Software instruction: A hybrid course is one that contains both online and classroom instruction. In a study performed at Middle Georgia State College, a software package was used as the main source of instruction in a hybrid college algebra course. The software tracked the number of hours it took for each student to meet the objectives of the course. In a sample of 44 students, the mean number of hours was 80.5, with a standard deviation of 51.4. Part: 0/2 Part 1 of 2 (a) Construct a 99.9% confidence interval for the mean number of hours it takes for a student to meet course objectives. Round the answers to one decimal place A 99.9% confidence interval for the mean number of hours it takes for a student to meet the course objectives is | | < μ< course objectives is
Answer #1

Solution :

Given that,

\bar x = 80.5

s = 51.4

n = 44

Degrees of freedom = df = n - 1 = 44 - 1 = 43

At 99.9% confidence level the t is ,

\alpha = 1 - 99.9% = 1 - 0.999 = 0.001

\alpha / 2 = 0.001 / 2 = 0.0005

t\alpha /2,df = t0.0005,43 = 3.532

Margin of error = E = t\alpha/2,df * (s /\sqrtn)

= 3.532 * ( 51.4/ \sqrt44)

= 27.4

The 99.9% confidence interval estimate of the population mean is,

\bar x - E < \mu < \bar x + E

80.5 - 27.4 < \mu < 80.5 + 27.4

53.1 < \mu < 107.9

(53.1,107.9)